# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ Abstract
Recent advances in object detectors have led to their adoption for industrial uses. However, their deployment in critical applications is hindered by the inherent lack of reliability of neural networks and the complex structure of object detection models. To address these challenges, we turn to Conformal Prediction, a post-hoc procedure which offers statistical guarantees that are valid for any dataset size, without requiring prior knowledge on the model or data distribution. Our contribution is manifold: first, we formally define the problem of Conformal Object Detection (COD) and introduce a novel method, Sequential Conformal Risk Control (SeqCRC), that extends the statistical guarantees of Conformal Risk Control (CRC) to two sequential tasks with two parameters, as required in the COD setting. Then, we propose loss functions and prediction sets suited to applying CRC to different applications and certification requirements. Finally, we present a conformal toolkit, enabling replication and further exploration of our methods. Using this toolkit, we perform extensive experiments, yielding a benchmark that validates the investigated methods and emphasizes trade-offs and other practical consequences.
Index Terms:
Conformal Prediction, Conformal Risk Control, Object Detection, AI Safety, Uncertainty Quantification
Figure 1: Conformal Object Detection Pipeline. Our method corrects the output of a given object detector $f$ with finite-sample probabilistic guarantees at a desired error rate $\alpha$. The figure illustrates the key stages: (1) Red bounding boxes show the initial raw predictions from the object detector $f(X_{i})$ which are then filtered by Non-Maximum Suppression to a smaller set. (2) Conformal Confidence Thresholding applies a statistically grounded confidence threshold, resulting in selected predictions shown with yellow bounding boxes with a guaranteed risk. (3) Conformal Localization corrects the previous dashed yellow boxes to build the final statistically valid purple bounding boxes. For each localized object, Conformal Classification constructs a prediction set of labels (e.g., "Set: {Boat, Plane}") from the model’s predictions, guaranteed to contain the true class with high probability. These steps together control a user-defined risk in localization and classification at a predetermined level $\alpha^{\mathrm{tot}}$.
@@ I Introduction
Object Detection (OD) has emerged as a pivotal task in computer vision, aiming to localize and classify objects within images or video sequences. It has had a remarkable impact in diverse fields, ranging from autonomous vehicles [1] to medical imaging [2]. However, there is an increasing demand to certify predictions of OD models within critical systems [3, 4, 5, 6].
In this light, while existing Uncertainty Quantification (UQ) methods are generally Bayesian [7, 8, 9] or heuristic [10, 11, 12], our approach emphasizes distribution-free UQ methods [13]. These methods neither assume prior knowledge of the data distribution nor are restricted to a particular class of predictive models. They are statistically valid even with a finite number of examples (i.e. non-asymptotic). When combining object detection with these methodologies, most notably Conformal Prediction (CP), we take an a posteriori strategy. This post-hoc approach is theoretically sound and computationally lightweight, allowing for rapid integration into real-world systems. Moreover, while the existing literature on conformal prediction for OD [14, 15, 16] and tolerance regions for OD [17, 18] focuses on specific steps of the whole OD pipeline or is sometimes tailored to specific models or tasks, our approach addresses a more unifying framework. It is applicable to virtually any object detectors, and readily deployable in existing systems, ensuring reliability without a comprehensive overhaul of established frameworks, and at very little cost.
@@
    In Section V, we empirically validate the correctness, efficiency, and broad applicability of our framework, via the first large benchmark of Conformal OD in the literature.
@@
    Finally, we provide a comprehensive, open-source codebase111 https://github.com/leoandeol/cods—a repository of models and conformal procedures—that facilitates the development of new guarantees, loss functions, and methodologies.
@@ I-A Related Works
Conformal prediction for OD. Conformal prediction [13, 19] and variants [20, 21, 22] are natural post-processing methods that offer statistically valid uncertainty estimates. Some applications to object detection have been proposed recently, notably through conformal prediction in [14, 16, 15, 23] and tolerance regions in [17, 18]. These two frameworks build and calibrate post-hoc prediction sets with guarantees, although slightly different ones. Most works have focused, however, only on the localization task (i.e., a regression task on the coordinates of a bounding box on the image). Two exceptions are [23], and [17]. The former, includes a guarantee on the classification task, while also producing conformal corrections in localization conditional on the class of the predicted object. The latter, builds all-encompassing sets, including confidence, classification and localization that are specifically tailored to Faster R-CNN models. However, none offers a holistic framework to build sets. That is, prediction sets that would be valid for virtually all families of OD models with simultaneous statistical guarantees on the number of object, as well as their localization, and classification.
A similar theoretical approach to ours, with two parameters to tune sequentially, was recently and independently introduced in [24] for rank retrieval. However, their theoretical guarantee is either asymptotic or requires two data splits for finite samples, whereas our method provides a finite-sample guarantee using a single data split only. The framework of LTT [25] provides another method to build prediction sets, which can depend on more than one parameter, using a multiple testing approach. This framework has recently been applied to Rank Retrieval [24] and to Language Modeling [26]. It has also been applied to OD to simultaneously control precision and recall in a medical setting in [18].
Other approaches to UQ for OD. Leaving statistical guarantees aside, let us mention other useful methods to compute uncertainty estimates for OD. First, several heuristic approaches exist. MetaDetect [10] trains an auxiliary network on top of the object detector to estimate predictive uncertainty from handpicked metrics. It does so by learning to discriminate false positives from true positives (in terms of Intersection over Union, IoU), and to predict IoU. The method of [27] allows to calibrate the confidence score of predicted bounded boxes, in a position-aware manner. The resulting confidence of bounding boxes in all regions of the image is then improved in terms of calibration error. Estimating uncertainty from the gradient of the loss (with regards to the the weights) has also been studied in [28] and a comparison of the two approaches is discussed in [29]. Notably, both approaches offer significantly more insights into a bounding box’s likelihood of containing an object than the native confidence score.
@@ II Setting: Object Detection with Uncertainty Quantification
Both Object Detection (OD) and Uncertainty Quantification (UQ) for OD have long been dominated by heuristic methods. As we aim to bring guaranteed uncertainty quantification to this field, we formally introduce the setting, the problem considered, and previous works. Our new SeqCRC method and theoretical analysis are presented in Section III, while a large selection of OD-specific modeling choices are discussed in Section IV.
@@ II-A Problem Setup (without UQ)
The task of object detection is complex, as it involves both localizing and classifying, in an image, an a priori unknown number of objects. To that end, most models (e.g. [38, 39, 40, 41]) predict a large amount of bounding boxes with an additional confidence component. This score is further used to filter out most predicted boxes using a threshold to keep the most relevant bounding boxes. While existing models compute this score in different ways, and some may use additional scores [42], we only consider a single score, irrespective of its construction, to provide a holistic setting and method.
Formal Problem Setup. We denote by $\mathcal{X}$ the input (image) space. Each object on an image is modeled as an element of ×$\mathcal{B}\times\mathcal{C}$, where =+$\mathcal{B}=\mathbb{R}^{4}_{+}$ is the bounding box space and $\mathcal{C}=\left\{1,\cdots,K\right\}$ is the label space ($K$ is the number of labels). Each location vector $b=(b^{\leftarrow},b^{\uparrow},b^{\rightarrow},b^{\downarrow})\in\mathcal{B}$ contains the coordinates of the top-left and bottom-right corners of its corresponding bounding box, respectively $(b^{\leftarrow},b^{\uparrow})$ and $(b^{\rightarrow},b^{\downarrow})$. We assume that $b^{\leftarrow}\leq b^{\rightarrow}$ and $b^{\uparrow}\leq b^{\downarrow}$. [^2] The horizontal axis points to the right while the vertical axis heads downwards.
@@
Commonly in OD, the output for a predicted object is not a single label, but an estimated probability distribution on $\mathcal{C}=\{1,\cdots,K\}$, such as a softmax output. Further recall that we work with a predictor $f$ that outputs a large, fixed number $N^{\text{pred}}$ of objects. Those objects are each composed of a bounding box, with an additional classification score (softmax output) and a confidence score in $[0,1]$. Such predictors are therefore functions $f:\mathcal{X}\rightarrow\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\text{pred}}}$, where $\Sigma_{K-1}$ denotes the $(K-1)$-simplex. [^3] That is, $\Sigma_{K-1}=\{c\in[0,1]^{K}:\sum_{i=1}^{K}c_{i}=1$}.
Importantly, a recurring post-processing step for object detectors is Non-Maximum Suppression (NMS), which consists in removing redundancies and thus significantly lowering the amount of predicted boxes. Its output lies in the reduced space (×$\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\mathrm{nms}}_{i}}$ where $N^{\mathrm{nms}}_{i}\ll N^{\text{pred}}$ and varies for each image $X_{i}$.
@@ III Conformal Object Detection
In this paper, we study an extension of Split Conformal Prediction called Conformal Risk Control (CRC) [53] and that adds flexibility for the user—a necessary feature for many critical applications. Next we start with some useful background on CRC, which only applies to prediction sets indexed by a single (scalar) parameter. Then, in Section III-B, we design and analyze an extension of CRC suited to our Object Detection setting, with three dependent scalar parameters.
@@ III-A Background: Conformal Risk Control
In this section we recall how conformal risk control proceeds in the post-hoc setting, [^4] By post-hoc we refer to a procedure conducted on top of a pretrained model’s predictions. for a general supervised learning task such as classification or regression. Assume that we are given a deterministic predictor $f$ [^5] The predictor is in fact required to be independent from the test and the calibration data. Roughly speaking, this corresponds to a setting where the model $f$ is trained and validated on a different split or pretrained. To construct a prediction set, we require a calibration dataset $\mathcal{D}_{\text{cal}}=(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ of $n$ examples and a test point $X_{\mathrm{test}}\in\mathcal{X}$. We then build a prediction set $\Gamma_{\lambda}(X_{\mathrm{test}})$666 Equivalently, prediction sets can be expressed in terms of a non-conformity score. See [54] for details. with the goal of controlling a loss $\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$ below a user-defined threshold $\alpha$. For example, in multiclass classification where $\mathcal{Y}=\{1,\ldots,K\}$, we can work with $\Gamma_{\lambda}(X_{\mathrm{test}})\subset\mathcal{Y}$ and $\ell(y,S)=\mathbbm{1}_{y\notin S}$. Multiple other examples exist, with non-binary losses or prediction sets $\Gamma_{\lambda}(X_{\mathrm{test}})$ that are not necessarily subsets of $\mathcal{Y}$ as, e.g., for the localization task (in that case, the prediction set is the predicted box after application of a margin, see Section IV-C). For notational simplicity, we denote the random losses777 These are referred to as random losses because they are functions of random variables, hence random variables themselves. by $L_{i}(\lambda)=\ell(Y_{i},\Gamma_{\lambda}(X_{i}))$, for $i=1,...,n$. We also denote the random loss of the test sample by $L_{\mathrm{test}}(\lambda)=\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$.
@@
where the lower bound holds under an additional assumption of continuity on the losses.
@@ III-B Sequential Conformal Risk Control with 1+2 parameters
In the classical setting recalled in Section III-A, we control a loss on a prediction set indexed by a single parameter $\lambda$. Next, in order to match the object detection setting of Section II, we consider three parameters: $\lambda^{\mathrm{cnf}}$ for confidence thresholding, $\lambda^{\mathrm{loc}}$ for the localization task, and $\lambda^{\mathrm{cls}}$ for the classification task. The localization and classification tasks are dependent on the confidence task, and existing theory for conformal prediction cannot handle this situation. Therefore we introduce Sequential Conformal Risk Control (SeqCRC), a new method that extends the CRC algorithm to handle 1+2 sequentially dependent parameters, which is well suited for OD tasks.
Notation and assumptions. We denote by $L_{i}^{j}$, with $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$, the loss functions for the confidence ($L^{\mathrm{cnf}}$), localization ($L^{\mathrm{loc}}$) and classification ($L^{\mathrm{cls}}$) tasks. For each task $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$ we also write $\lambda^{j}\in\Lambda^{j}$ for the associated parameter and $\alpha^{j}$ for the desired error rate. We assume that the $\Lambda^{j}\subset\mathbb{R}$ are compact, and set $\bar{\lambda}^{j}=\max\Lambda^{j}$. (Our examples in Section IV are of the form $\Lambda^{j}=\bigl{[}0,\bar{\lambda}^{j}\bigr{]}$.)

### v2

@@ Abstract
Recent advances in object detectors have led to their adoption for industrial uses. However, their deployment in safety-critical applications is hindered by the inherent lack of reliability of neural networks and the complex structure of object detection models. To address these challenges, we turn to Conformal Prediction, a post-hoc predictive uncertainty quantification procedure with statistical guarantees that are valid for any dataset size, without requiring prior knowledge on the model or data distribution. Our contribution is manifold. First, we formally define the problem of Conformal Object Detection (COD). We introduce a novel method, Sequential Conformal Risk Control (SeqCRC), that extends the statistical guarantees of Conformal Risk Control to two sequential tasks with two parameters, as required in the COD setting. Then, we present old and new loss functions and prediction sets suited to applying SeqCRC to different cases and certification requirements. Finally, we present a conformal toolkit for replication and further exploration of our method. Using this toolkit, we perform extensive experiments that validate our approach and emphasize trade-offs and other practical consequences.
Figure 1: Conformal Object Detection Pipeline. Our method corrects the output of a given object detector $f$ with finite-sample probabilistic guarantees at a desired error rate $\alpha$. The figure illustrates the key stages: (1) Red bounding boxes show the initial raw predictions from the object detector $f(x)$ which are then filtered by Non-Maximum Suppression to a smaller set. (2) Conformal Confidence Thresholding applies a statistically grounded confidence threshold, resulting in selected predictions shown with yellow bounding boxes with a guaranteed risk. (3) Conformal Localization corrects the previous dashed yellow boxes to build the final statistically valid purple bounding boxes. For each localized object, Conformal Classification constructs a prediction set of labels (e.g., "Set: {Boat, Plane}") from the model’s predictions, guaranteed to contain the true class with high probability. These steps together control a user-defined risk in localization and classification at a predetermined level $\alpha^{\mathrm{tot}}$.
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Object Detection (OD) has emerged as a pivotal task in computer vision, which aims at localizing and classifying objects within images or video sequences. It has had a remarkable impact in diverse fields, ranging from autonomous vehicles [1] to medical imaging [2]. However, there is an increasing demand to certify predictions of OD models within safety-critical systems [3, 4, 5, 6].
In this light, while most Uncertainty Quantification (UQ) methods are Bayesian [7, 8, 9] or heuristic [10, 11, 12], our approach emphasizes distribution-free UQ methods [13]. These methods neither assume prior knowledge of the data distribution nor are restricted to a particular class of predictive models. They are statistically valid even with a finite number of examples (i.e. non-asymptotic). When combining object detection with these methodologies, most notably Conformal Prediction (CP), we take an a posteriori strategy. This post-hoc approach is theoretically sound and computationally lightweight, allowing for rapid integration into real-world systems. Existing literature on conformal prediction for OD [14, 15, 16] and on tolerance regions for OD [17, 18] typically targets specific steps of the OD pipeline or individual model families. In contrast, our framework covers the entire detection pipeline within a single unified approach that applies across models and tasks, while providing statistical guarantees. It is applicable to virtually any object detectors, and readily deployable in existing systems, ensuring reliability without a comprehensive overhaul of established frameworks, and at very little cost.
@@
    In Section V, we empirically validate the correctness, efficiency, and broad applicability of our framework, via the first extensive experiments of Conformal OD in the literature.
@@
    Finally, we provide a comprehensive, open-source codebasehttps://github.com/leoandeol/cods—a repository of models and conformal procedures—that facilitates the development of new guarantees, loss functions, and methodologies.
In the supplemental material, we provide a proof of our theoretical guarantee, more detailed pseudocodes, as well as additional remarks on modeling choices and experiments.
@@ I-A Related Works
Conformal prediction for OD. Conformal prediction [13, 19] and variants [20, 21, 22] are natural post-processing methods that offer statistically valid uncertainty estimates. Some applications to object detection have been proposed recently, notably through conformal prediction and tolerance regions. These two frameworks build and calibrate post-hoc prediction sets with statistical guarantees, although slightly different ones. To the best of our knowledge, most such applications to OD, e.g., [14, 16, 15, 23], only consider the localization task, i.e., predicting the coordinates of bounding boxes around objects on an image. Two exceptions are [24] and [17]. The former provides a guarantee on the classification task, while also producing conformal corrections for localization that are conditional on the class of the predicted object. The other work [17] builds all-encompassing sets (including confidence, classification and localization) that are specifically tailored to Faster R-CNN models. However, to the best of our knowledge, no work offers a holistic framework to build sets. That is, prediction sets that would be valid for virtually all families of OD models, with statistical guarantees on both the number of predicted objects and their localization and classification accuracies.
A similar approach to ours, with two parameters to tune sequentially, was recently and independently introduced in [25] for rank retrieval. However, their theoretical guarantee is either asymptotic or requires two data splits for finite samples, whereas our method provides a finite-sample guarantee using a single data split only. The framework of LTT [22] provides another method to build prediction sets, which can depend on more than one parameter, using a multiple testing approach. This framework has recently been applied to language modeling [26] as well as to rank retrieval [25]. It has also been applied to OD for medical data [18], for the problem of simultaneously controlling precision and recall.
Other approaches to UQ for OD. Leaving statistical guarantees aside, let us mention other useful methods to compute uncertainty estimates for OD. First, several heuristic approaches exist. MetaDetect [10] trains an auxiliary network on top of the object detector to estimate predictive uncertainty from handpicked metrics. It does so by learning to discriminate false positives from true positives (in terms of Intersection over Union, IoU), and to predict IoU. The method of [27] allows to calibrate the confidence score of predicted bounded boxes, in a position-aware manner. The resulting confidence of bounding boxes in all regions of the image is then improved in terms of calibration error. Estimating uncertainty from the gradient of the loss (with regards to the weights) has also been studied in [28] and a comparison of the two approaches is discussed in [29]. Notably, both approaches offer significantly more insights into a bounding box’s likelihood of containing an object than the native confidence score.
@@ II Setting: Object Detection with Uncertainty Quantification
Both Object Detection (OD) and Uncertainty Quantification (UQ) for OD have long been dominated by heuristic methods. As we aim to bring guaranteed uncertainty quantification to this field, we start by formally introducing the setting. Our new SeqCRC method and theoretical analysis are presented in Section III, while a large selection of OD-specific modeling choices are discussed in Section IV.
@@ II-A Problem Setup (without UQ)
Object detection on images is a complex task, as it involves both localizing and classifying an a priori unknown number of objects for each input image. To that end, most models (e.g. [38, 39, 40, 41]) predict a large amount of bounding boxes with an additional confidence component. This score is further used to filter out most predicted boxes using a threshold to keep the most relevant bounding boxes. While existing models compute this score in different ways, and some may use additional scores [42], we only consider a single score, irrespective of its construction, to provide a holistic setting and method.
Formal Problem Setup. We denote by $\mathcal{X}$ the input (image) space. Each object on an image is modeled as an element of $\mathcal{B}\times\mathcal{C}$, where $\mathcal{B}=\mathbb{R}^{4}_{+}$ is the bounding box space and $\mathcal{C}=\left\{1,\cdots,K\right\}$ is the label space ($K$ is the number of labels). Each location vector $b=(b^{\leftarrow},b^{\uparrow},b^{\rightarrow},b^{\downarrow})\in\mathcal{B}$ contains the coordinates of the top-left and bottom-right corners of its corresponding bounding box, respectively $(b^{\leftarrow},b^{\uparrow})$ and $(b^{\rightarrow},b^{\downarrow})$. We assume that $b^{\leftarrow}\leq b^{\rightarrow}$ and $b^{\uparrow}\leq b^{\downarrow}$.The horizontal axis points to the right while the vertical axis heads downwards.
@@
Commonly in OD, the output for a predicted object is not a single label, but an estimated probability distribution on $\mathcal{C}=\{1,\cdots,K\}$, such as a softmax output. Further recall that we work with a predictor $f$ that outputs a large, fixed number $N^{\text{pred}}$ of objects. Each of these objects is composed of a bounding box, with an additional classification score (softmax output) and a confidence score in $[0,1]$. Such predictors are therefore functions $f:\mathcal{X}\rightarrow\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\text{pred}}}$, where $\Sigma_{K-1}$ denotes the (K1)(K-1)-simplex.That is, $\Sigma_{K-1}=\{c\in[0,1]^{K}:\sum_{i=1}^{K}c_{i}=1$}.
Importantly, a recurring post-processing step for object detectors is Non-Maximum Suppression (NMS), which consists in removing redundancies and thus significantly lowering the amount of predicted boxes. Its output lies in the reduced space $\left(\mathcal{B}\times\Sigma_{K-1}\times[0,1]\right)^{N^{\mathrm{nms}}}$ where the number $N^{\mathrm{nms}}$ depends on $x$ and is typically much smaller than $N^{\text{pred}}$.
Similarly, each image $x$ is assigned a true complex label $y\in(\mathcal{B}\times\mathcal{C})^{|y|}$ given by the list (or sequence) of all annotated ground-truth objects on $x$ (we write $|y|$ for the number of objects). We will sometimes use $\mathcal{Y}$ to denote the set of all such complex labels $y$.
@@ III Conformal Object Detection
In this paper, we study an extension of Split Conformal Prediction called Conformal Risk Control (CRC) [53] and that adds flexibility for the user—a necessary feature for many applications. Next we start with some useful background on CRC, which only applies to prediction sets indexed by a single (scalar) parameter. Then, in Section III-B, we design and analyze an extension of CRC suited to our Object Detection setting, with three dependent scalar parameters.
@@ III-A Background: Conformal Risk Control
In this section we recall how conformal risk control proceeds in the post-hoc setting,By post-hoc we refer to a procedure conducted on top of a pretrained model’s predictions. for a general supervised learning task such as classification or regression (we momentarily work with generic input and label spaces $\mathcal{X}$ and $\mathcal{Y}$). Assume that we are given a deterministic predictor $f$.The predictor is in fact required to be independent from the test and the calibration data. Roughly speaking, this corresponds to a setting where the model $f$ is trained and validated on a different split or pretrained. To construct a prediction set, we require a calibration dataset $\mathcal{D}_{\text{cal}}=(X_{i},Y_{i})_{i=1}^{n}\in(\mathcal{X}\times\mathcal{Y})^{n}$ of $n$ examples and a test point $X_{\mathrm{test}}\in\mathcal{X}$. We then build a prediction set $\Gamma_{\lambda}(X_{\mathrm{test}})$Equivalently, prediction sets can be expressed in terms of a non-conformity score. See [54] for details. with the goal of controlling a loss $\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$ below a user-defined threshold $\alpha$. For example, in multiclass classification where $\mathcal{Y}=\{1,\ldots,K\}$, we can work with $\Gamma_{\lambda}(X_{\mathrm{test}})\subset\mathcal{Y}$ and $\ell(y,S)=\mathbbm{1}_{y\notin S}$. Multiple other examples exist, with non-binary losses or with prediction sets $\Gamma_{\lambda}(X_{\mathrm{test}})$ that are not necessarily subsets of $\mathcal{Y}$ as, e.g., for the localization task (in that case, the prediction set is the predicted box after application of a margin, see Section IV-C). For notational simplicity, we denote the random lossesThese are referred to as random losses because they are functions of random variables, hence random variables themselves. by $L_{i}(\lambda)=\ell(Y_{i},\Gamma_{\lambda}(X_{i}))$, for i=1,,ni=1,...,n. We also denote the random loss of the test sample by $L_{\mathrm{test}}(\lambda)=\ell(Y_{\mathrm{test}},\Gamma_{\lambda}(X_{\mathrm{test}}))$.
@@
where the lower bound holds under an additional continuity-type assumption on the random losses.
@@ III-B Sequential Conformal Risk Control with 1+2 parameters
In the classical setting recalled in Section III-A, one can control a loss on a prediction set indexed by a single parameter $\lambda$. In all the sequel, we work within the object detection setting of Section II. The calibration and test data correspond to random pairs $(X_{i},Y_{i})$, $i\in\{1,\ldots,n\}\cup\{\text{test}\}$, where $X_{i}\in\mathcal{X}$ is an image and $Y_{i}\in\mathcal{Y}$ denotes its true list of bounding boxes and classes. More importantly, we now consider three parameters: $\lambda^{\mathrm{cnf}}$ for confidence thresholding, $\lambda^{\mathrm{loc}}$ for the localization task, and $\lambda^{\mathrm{cls}}$ for the classification task. The localization and classification tasks are dependent on the confidence task, and existing theory for conformal prediction cannot handle this situation. Therefore we introduce Sequential Conformal Risk Control (SeqCRC), a new method that extends the CRC algorithm to handle 1+2 sequentially dependent parameters, which is well suited for OD tasks.
III-B1 Notation and assumptions
We denote by $L_{i}^{j}$, with $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$, the loss functions for the confidence ($L^{\mathrm{cnf}}$), localization ($L^{\mathrm{loc}}$) and classification ($L^{\mathrm{cls}}$) tasks. For each task $j\in\{\mathrm{cnf},\mathrm{loc},\mathrm{cls}\}$ we also write $\lambda^{j}\in\Lambda^{j}$ for the associated parameter and $\alpha^{j}$ for the desired error rate. We assume that the $\Lambda^{j}\subset\mathbb{R}$ are compact, and set $\bar{\lambda}^{j}=\max\Lambda^{j}$. (Our examples in Section IV are of the form $\Lambda^{j}=\bigl[0,\bar{\lambda}^{j}\bigr]$.)